OpenStudy (anonymous):

Find the midpoint between (8,-13) and (1,-7).

OpenStudy (anonymous):

take the average of each coordinate

OpenStudy (anonymous):

Label your coordinate points (x1,y1) and (x2,y2) and and use the midpoint formula \[(x _{1}+x _{2}/2 , y _{1}+y _{2}/2)\]

OpenStudy (anonymous):

get \[(\frac{9}{2},-10)\]

OpenStudy (anonymous):

hello again! you are still here?

OpenStudy (anonymous):

just add up and divide by 2. easy enough you may could do it in your head

OpenStudy (anonymous):

More homework... 17 more problems, to be exact. (: & im lost... what am i adding and dividing? Honestly, i have no clue what a midpoint is....

OpenStudy (anonymous):

midpoint means point half way between the other two

OpenStudy (anonymous):

Welll that would make sense...

OpenStudy (anonymous):

so you compute as follows: \[(\frac{8+1}{2},\frac{-13-7}{2})\]

OpenStudy (anonymous):

hence the answer i sent

OpenStudy (anonymous):

17 more to go

OpenStudy (anonymous):

thats just plugging it in to the formula, right? THen i gotta solve, then ill get the answer you wrote?

OpenStudy (anonymous):

yes that is it. add up and divide by two. one of the easier math problems

OpenStudy (anonymous):

you realize the answer is a point so it should look like \[(a,b)\]

OpenStudy (anonymous):

right, okay. Well hold on, i gotta write it all out and such.